If the velocity of an object is doubled and its mass is halved, which of the following statements is true about the object? Select one: a. Its momentum is halved; its kinetic energy doubles. b. Its momentum stays the same; its kinetic energy doubles. c. Its momentum is halved; its kinetic energy quadruples. d. Its momentum is halved; its kinetic energy is also halved.
Answer #1

Option b is correct.

Its momentum stays the same; its kinetic energy doubles.

Kinetic Energy,K = (1/2)*m*v^2

Momentum,P = m*v

If velocity is doubled and mass is halved,

Kinetic Energy = (1/2)*(m/2)*(2v)^2 = 2K

Momentum = (m/2)*(2v) = P